# CytoNet: A Foundation Model for the Human Cerebral Cortex

### v1

@@ 2 Discussion
We demonstrated that CytoNet learns cortical cytoarchitecture from large-scale histological data without requiring manual annotations. The model encodes laminar and areal organization into a feature space that generalizes across subjects, aligns with cortical morphology, and supports diverse applications including area classification, layer segmentation, and exploratory parcellation. These results show that anatomical proximity provides an effective training signal, turning spatial continuity into a self-supervised objective that captures the latent factors of cortical organization. Together, they establish CytoNet as a biologically grounded and practically scalable foundation for systematic analysis of cortical histology at unprecedented scale, capable of integrating cytoarchitectonic patterns across brains while preserving both global consistency and individual variability.
@@
The properties of these representations can be traced back to the proposed SpatialNCE training strategy. SpatialNCE derives its training signal by assuming that spatially close samples are often semantically consistent. In terms of cortical organization, this translates to prior knowledge that textures observed in proximity often show architectural variations of the same functional modules. This principle distinguishes CytoNet from other self-supervised learning approaches, both contrastive [118, 123, 117, 122] and non-contrastive [121, 127]. It directly addresses the limitations we encountered with the SimCLR approach [118], which sometimes performed even below scratch training because it relied on augmentation-invariant confounding cues such as vascular patterns or local curvature rather than cytoarchitecture (shortcut learning, [120]). Why, then, does augmentation-based self-supervised learning like SimCLR succeed in natural images but fail here? As argued by [130], augmentations implicitly define which features are treated as task-relevant content and which as task-irrelevant style. In our setting, standard augmentations (e.g., intensity distortions, noise, geometric transforms) leave confounds like vascular patterns or folding geometry intact, encouraging the model to treat them as content. This limitation reflects both the dataset and the chosen patch size: While choosing sufficiently large fields of view to capture laminar profiles inevitably includes confounds, smaller inputs could remove confounds but sacrifice important context. On the other hand, stronger augmentations (e.g., elastic deformations) may lead to undesired distortions of cytoarchitectonic structure. From this point of view, proximity-based similarity measures serve as a biologically grounded augmentation, leveraging cortical continuity across 3D space and across subjects to generate natural variations that disrupt confounds while preserving cytoarchitecture. Beyond cytoarchitecture, SpatialNCE provides an intuitive and easy-to-use framework for multimodal self-supervision: any dataset anchored into a common spatial frame can be integrated under the same loss without modification, with only modest requirements for spatial alignment. Although demonstrated here for cortical cytoarchitecture, the principle generalizes to other domains where spatial proximity systematically relates to semantic structure (e.g., remote sensing), establishing it as a unifying paradigm for multimodal representation learning.
@@
CytoNet’s training objective defines similarity based on the Euclidean distance between image patches. However, spatial coordinates are never provided as input to the model. Keeping this in mind, the organization of the learned feature space yields two key insights. First, the strong alignment between the learned feature space and anatomical location shows that local image information reliably reflects the spatial continuity of cortical organization. Second, the objective is at the same time flexible enough to deviate from this imposed geometry when needed: although its objective promotes similarity across brains, CytoNet recovers distinct brain-specific clusters with consistent internal organization. A striking example is the separation along the central sulcus: despite their adjacency in 3D space, the motor and somatosensory areas that are located anterior and posterior to the central sulcus form distinct regions in the feature space, illustrating how CytoNet recovers sharp cytoarchitectonic contrasts.
@@
CytoNet provides a versatile foundation for automated brain mapping tasks, supporting applications such as brain area classification, cortical layer segmentation, and data-driven area discovery. Unlike earlier methods [111, 132] it generalizes across entire brains and multiple individuals without task-specific retraining, bringing fully automated mapping within reach. In practice, CytoNet outputs still require post-processing (e.g., spatial smoothing and topological constraints), but the model shifts the bottleneck from manual annotation to scalable, data-driven analysis, making it feasible to process entire brains, compare across individuals, and handle terabyte- to petabyte-scale datasets. To the best of our knowledge, CytoNet currently represents the most precise method for cytoarchitectonic area classification, particularly when applied to brains without annotated training data.
@@
In summary, CytoNet represents a step towards a new generation of approaches to study brain organization that are: (i) anatomically rooted, capturing fundamental structural principles, (ii) scalable, enabling dense and reproducible mapping across whole brains, (iii) general, applicable across regions, subjects, and imaging modalities, and (iv) extensible, providing a foundation for multi-modal integration and holistic models of brain organization.
@@ 3.4 Self-supervised pretraining of CytoNet
Data augmentation was applied to capture typical variations in the data, following the data augmentation strategy detailed in [132], which is briefly summarized below. Image patches were randomly rotated by $\theta\in U\left[-\pi,+\pi\right]$ ($U[a,b]$: uniform distribution over $[a,b]$), patch center positions were translated in a random direction by $d\sim U\left[0\textrm{mm},0.2\textrm{mm}\right]$, and mirrored vertically with a probability of $50\%$. Pixel intensities $x\in\left[0,1\right]$ were randomly augmented using unbiased gamma augmentation [107] $\alpha x^{\gamma}+\beta$ with parameters $\alpha\in U[0.9,1.0]$, $\beta\in U[-0.1,+0.1]$, $\gamma=\frac{\log\left(0.5+2^{-0.5Z}\right)}{\log\left(0.5-2^{-0.5Z}\right)}$, $Z\sim U[-0.05,+0.05]$. In addition, images were blurred with an isotropic Gaussian $G_{\sigma}(x)$ of kernel size $\sigma\sim U[0.125,1.0]$, or sharpened according to $x+\delta(G_{\sigma_{u}}(x)-x)$ with ${\sigma}_{u}\sim U[0.125,1.0]$, $\delta\sim U[0.5,1.5]$ with probability $25\%$, respectively.
@@ 3.7 Predicting structural variations in cytoarchitecture
Extraction of morphological features from BigBrain
@@
Extraction of layer-wise cell-density from microscopic image patches
@@
BigBrain intensity profiles
@@
Correlative analysis of CytoNet features and Bigbrain intensity profiles
To assess the information encoded in CytoNet representations, we conducted a correlative analysis between feature embeddings and known morphological and structural properties of the cerebral cortex. Learned feature representations of CytoNet-ViT (1M) were compared with structural features extracted from brain B2020, including cortical curvature, cortical thickness, layer-specific thickness, cutting direction, and layer-wise cell densities. In parallel, we conducted the same analysis using depth-wise intensity profiles (see above) to serve as a baseline for comparison.
@@
3.8 Data-driven parcellation of cortical areas
Dataset
@@ 3.9 Computational setup
3.10 Use of Large-Language Models
ChatGPT (version 5) was used to prepare the manuscript to improve the clarity and brevity of the text. All generated text was revised, checked for correctness, and accepted by all authors.
@@ 4 Acknowledgement
This project received funding from the European Union’s Horizon 2020 Research and Innovation Programme, grant agreement 101147319 (EBRAINS 2.0 Project), the Helmholtz Association port-folio theme “Supercomputing and Modeling for the Human Brain”, the Helmholtz Association’s Initiative and Networking Fund through the Helmholtz International BigBrain Analytics and Learning Laboratory (HIBALL) under the Helmholtz International Lab grant agreement InterLabs-0015, from HELMHOLTZ IMAGING, a platform of the Helmholtz Information & Data Science Incubator [X-BRAIN, grant number: ZT-I-PF-4-061], and from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the National Research Data Infrastructure – NFDI 46/1 – 501864659. Computing time was granted through JARA on the supercomputer JURECA-DC at Jülich Supercomputing Centre (JSC). The authors declare no competing interests.

### v2

@@ 3 Discussion
We demonstrated that CytoNet learns cortical cytoarchitecture from large-scale histological data without requiring manual annotations. The model encodes laminar and areal organization into a feature space that generalizes across subjects, aligns with cortical morphology, and supports diverse applications including area classification, layer segmentation, structure prediction, and exploratory parcellation. In addition, CytoNet features decode widely used functional network parcellations, supporting a direct link between cellular-scale microarchitecture and macroscale functional systems. These results show that anatomical proximity provides an effective training signal, turning spatial continuity into a self-supervised objective that captures the latent factors of cortical organization. Together, they establish CytoNet as a biologically grounded and practically scalable foundation for systematic analysis of cortical histology at whole-brain scale, capable of integrating cytoarchitectonic patterns across brains while preserving both global consistency and individual variability.
@@
The properties of the learned feature space can be traced back to the proposed SpatialNCE training strategy. SpatialNCE derives its training signal by assuming that spatially close samples are often semantically consistent. In terms of cortical organization, this translates to prior knowledge that textures observed in proximity often show architectural variations of the same functional modules. This principle distinguishes CytoNet from other self-supervised learning approaches, both contrastive [19, 47, 16, 39] and non-contrastive [37, 20]. It directly addresses the limitations we encountered with the SimCLR approach [19], which sometimes performed even below scratch training because it relied on augmentation-invariant confounding cues such as vascular patterns or local curvature rather than cytoarchitecture (shortcut learning, [33]). Why, then, does augmentation-based self-supervised learning like SimCLR succeed in natural images but fail here? As argued by [49], augmentations implicitly define which features are treated as task-relevant content and which as task-irrelevant style. In our setting, standard augmentations (e.g., intensity distortions, noise, geometric transforms) leave confounds like vascular patterns or folding geometry intact, encouraging the model to treat them as content. This limitation reflects both the dataset and the chosen patch size: While choosing sufficiently large fields of view to capture laminar profiles inevitably includes confounds, smaller inputs could remove confounds but sacrifice important context. On the other hand, stronger augmentations (e.g., elastic deformations) may lead to undesired distortions of cytoarchitectonic structure. From this point of view, proximity-based similarity measures serve as a biologically grounded augmentation, leveraging cortical continuity across 3D space and across subjects to generate natural variations that disrupt confounds while preserving cytoarchitecture. Beyond cytoarchitecture, SpatialNCE provides an intuitive and easy-to-use framework for multimodal self-supervision: any dataset anchored into a common spatial frame can be integrated under the same loss without modification, with only modest requirements for spatial alignment. Although demonstrated here for cortical cytoarchitecture, the principle generalizes to other domains where spatial proximity systematically relates to semantic structure (e.g., remote sensing), making it an attractive unifying paradigm for multimodal representation learning.
@@
CytoNet’s training objective defines similarity based on the Euclidean distance between sampling locations of image patches. However, spatial coordinates are never provided as input to the model. Keeping this in mind, the organization of the learned feature space yields two key insights: First, the strong alignment between the learned feature space and anatomical location shows that local image information reliably reflects the spatial continuity of cortical organization. Second, the objective is at the same time flexible enough to deviate from this imposed geometry when needed: although its objective promotes similarity across brains, CytoNet recovers distinct brain-specific clusters with consistent internal organization. A striking example is the separation along the central sulcus: despite their adjacency in 3D space, the motor and somatosensory areas that are located anterior and posterior to the central sulcus form distinct regions in the feature space, illustrating how CytoNet recovers sharp cytoarchitectonic contrasts.
@@
CytoNet provides a versatile foundation for automated brain mapping tasks, supporting applications such as brain area classification, cortical layer segmentation, and data-driven parcellation. Unlike earlier methods [75, 72], it generalizes across entire brains and multiple individuals without task-specific retraining, bringing fully automated mapping within reach. In our experiments, CytoNet represented the most precise method for cytoarchitectonic area classification, particularly when applied to brains without annotated training data. In practice, CytoNet outputs still require post-processing (e.g., spatial smoothing and topological constraints), but the model shifts the bottleneck from manual annotation to scalable, data-driven analysis, making it feasible to process entire brains, compare across individuals, and handle whole-brain scale datasets.
@@
Beyond structural mapping, CytoNet features captured microarchitectural signatures of macroscale functional organization, as evidenced by high performance in predicting functional network parcellations. While functional systems exhibit strong spatial smoothness, CytoNet provided predictive information beyond what can be explained by spatial coordinates alone and exceeded established low-dimensional gradient baselines. The improved performance when combining CytoNet features with spatial coordinates suggests that cytoarchitecture and cortical position contribute complementary signals. This result supports the view that cellular architecture provides a substrate for systems-level functional differentiation [59, 41] and motivates future work integrating microarchitectural and neuroimaging foundation models in a common representational space.
In summary, CytoNet represents a step towards a new generation of approaches to study brain organization that are: (i) anatomically rooted, capturing fundamental structural principles, (ii) scalable, enabling dense and reproducible mapping across whole brains, (iii) general, applicable across regions, subjects, and imaging modalities, and (iv) extensible, providing a foundation for multimodal integration and holistic models of brain organization.
@@ 4.4 Self-supervised pretraining of CytoNet
Data augmentation was applied to capture typical variations in the data, following the data augmentation strategy detailed in [72], which is briefly summarized below. Image patches were randomly rotated by $\theta\sim U\left[-\pi,+\pi\right]$ ($U[a,b]$: uniform distribution over $[a,b]$), patch center positions were translated in a random direction by $d\sim U\left[0\textrm{mm},0.2\textrm{mm}\right]$, and mirrored vertically with a probability of $50\%$. Pixel intensities $x\in\left[0,1\right]$ were randomly augmented using unbiased gamma augmentation [68] $\alpha x^{\gamma}+\beta$ with parameters $\alpha\sim U[0.9,1.1]$, $\beta\sim U[-0.1,+0.1]$, $\gamma=\frac{\log\left(0.5+2^{-0.5}Z\right)}{\log\left(0.5-2^{-0.5}Z\right)}$, $Z\sim U[-0.05,+0.05]$. In addition, images were blurred with an isotropic Gaussian $G_{\sigma}(x)$ of kernel size $\sigma\sim U[0.125,1.0]$, or sharpened according to $x+\delta(G_{\sigma_{u}}(x)-x)$ with ${\sigma}_{u}\sim U[0.125,1.0]$, $\delta\sim U[0.5,1.5]$ with probability $25\%$, respectively. After augmentation, images were standardized to $\left[-1,1\right]$ before being passed into the model.
@@ 4.7 Predicting structural variations in cytoarchitecture
Extraction of morphological features from BigBrain.
@@
Extraction of layer-wise cell-density from microscopic image patches.
@@
BigBrain intensity profiles.
@@
Correlative analysis of CytoNet features and BigBrain intensity profiles.
We conducted a correlative analysis between feature embeddings and known morphological and structural properties of the cerebral cortex. Learned feature representations of CytoNet-ViT (1M) were compared with structural features extracted from brain B2020, including cortical curvature, cortical thickness, layer-specific thickness, cutting direction, and layer-wise cell densities. In parallel, we conducted the same analysis using depth-wise intensity profiles (see above) to serve as a baseline for comparison.
@@
4.8 Decoding functional organization from cytoarchitecture
We evaluated whether CytoNet representations can decode macroscale functional organization on the cortical surface. CytoNet-ViT (1M) features were computed for BigBrain B2020, assigned to the BigBrain midsurface by nearest-neighbor lookup in 3D coordinate space, and mapped to fs_LR surface space for all decoding analyses.
As functional targets, we used the Yeo7 and Yeo17 network parcellations [87] as well as the DiFuMo atlas with 64 functional modes [23]. Yeo labels were obtained in BigBrain space from the BigBrainWarp toolbox [64]. DiFuMo64 was fetched in MNI152 space using the siibra tool suite [25] and mapped to fs_LR (32k) surface space [81] with nearest-neighbor resampling using neuromaps [54].
All experiments were performed in fs_LR (32k) surface space after mapping all features and targets to this space. CytoNet feature maps were warped from BigBrain space to fs_LR using BigBrainWarp. For baseline comparisons, we additionally included histological gradients (Hist-G1 and Hist-G2) [65], microstructural gradients derived from qT1 (Micro-G1 and Micro-G2) [71], and functional gradients (Func-G1 and Func-G2) [53], obtained via BigBrainWarp and evaluated in the same fs_LR space. We also evaluated a spatial baseline using fs_LR vertex coordinates (x,y,z), and a combined feature set concatenating CytoNet PCA features with coordinates.
To obtain low-dimensional CytoNet variants, principal component analysis (PCA) was fitted once on CytoNet features pooled across both hemispheres in fs_LR space and subsequently applied to each hemisphere. We evaluated PCA projections with 1, 2, and 32 components.
Functional labels were predicted using multinomial logistic regression with class balancing, implemented in scikit-learn. All input features were standardized within each training fold (zero mean, unit variance) and the transformation was applied to the held-out fold.
Spatially grouped cross-validation.
To reduce score inflation from spatial autocorrelation, we used grouped 5-fold cross-validation with groups defined by spatial surface superblocks. Superblocks were generated by 3-fold explicit isotropic mesh decimation of the fs_LR surface to obtain a coarser set of vertices, assigning each original vertex to its nearest remeshed vertex, and then clustering block centroids into 100 parcels per hemisphere using $k$-means. All vertices belonging to the same superblock were kept in the same fold.
Spin-based spatial null.
To test whether performance gains over the coordinate baseline can be explained by spatial smoothness alone, we performed a spin-based spatial permutation test on the Yeo labels in fs_LR space. For each hemisphere, labels were rotated on the spherical fs_LR representation using random rotations and reassigned to vertices by nearest-neighbor matching on the sphere, implemented with neuromaps [54]. Rotations were performed independently per hemisphere to preserve hemispheric topology, and the same spatially grouped cross-validation procedure was applied after each spin. We assessed statistical significance of the observed improvement in macro-F1 between CytoNet PCA–32 and the coordinate-only baseline using a one-sided test, with $p=(1+\#\{\Delta_{\mathrm{spin}}\geq\Delta_{\mathrm{obs}}\})/(1+N_{\mathrm{spin}})$. When testing multiple parcellations, $p$-values were corrected using Holm–Bonferroni. Performance was quantified by macro-F1 and reported as mean $\pm$ standard deviation across spatially grouped cross-validation folds. To reduce runtime, spin tests were conducted in parallel using 1024 CPU cores across 4 compute nodes of JURECA-DC (runtime Yeo7: 18 minutes; Yeo17: 43 minutes; DiFuMo64: 160 minutes).
4.9 Data-driven mapping of cortical areas
Dataset.
@@ 5 Acknowledgement
This project received funding from the European Union’s Horizon 2020 Research and Innovation Programme, grant agreement 101147319 (EBRAINS 2.0 Project), the Helmholtz Association port-folio theme “Supercomputing and Modeling for the Human Brain”, the Helmholtz Association’s Initiative and Networking Fund through the Helmholtz International BigBrain Analytics and Learning Laboratory (HIBALL) under the Helmholtz International Lab grant agreement InterLabs-0015, from HELMHOLTZ IMAGING, a platform of the Helmholtz Information & Data Science Incubator [X-BRAIN, grant number: ZT-I-PF-4-061], and from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the National Research Data Infrastructure – NFDI 46/1 – 501864659. Computing time was granted through JARA on the supercomputer JURECA-DC at Jülich Supercomputing Centre (JSC).
6 Author contributions
CS: Conceptualization, Data curation, Formal analysis, Funding acquisition, Investigation, Methodology, Project administration, Software, Validation, Visualization, Writing – original draft. ZB: Methodology, Software, Writing – review & editing. JK: Methodology, Software, Validation, Writing – review & editing. JT: Data curation, Methodology, Software, Writing – review & editing. KB: Formal analysis, Investigation, Validation, Writing – review & editing. HS: Methodology, Investigation, Validation, Software, Writing – review & editing. KA: Conceptualization, Funding acquisition, Methodology, Project administration, Resources, Supervision, Validation, Writing – review & editing. TD: Conceptualization, Funding acquisition, Methodology, Project administration, Software, Supervision, Validation, Writing – review & editing.
7 Declaration of Interests
The authors declare no competing interests.
8 Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT (versions 5.0 and 5.1) in order to improve the clarity and brevity of the text. After using this tool, the author(s) reviewed and edited the content as needed and take full responsibility for the content of the published article.